# Security Weaknesses of Copilot Generated Code in GitHub

### v1

@@ Abstract.
Modern code generation tools use AI models, particularly Large Language Models (LLMs), to generate functional and complete code. While such tools are becoming popular and widely available for developers, using these tools is often accompanied by security challenges, leading to insecure code merging into the code base. Therefore, it is important to assess the quality of the generated code, especially in terms of its security. Researchers have recently explored various aspects of code generation tools, including security. However, many open questions about the security of the generated code require further investigation, especially the security issues of automatically generated code in the wild. To this end, we conducted an empirical study by analyzing the security weaknesses in code snippets generated by GitHub Copilot that are found as part of publicly available projects hosted on GitHub. The goal is to investigate the types of security issues and their scale in real-world scenarios (rather than crafted scenarios). To this end, we identified 435 code snippets generated by GitHub Copilot from publicly available projects. We then conducted extensive security analysis to identify Common Weakness Enumeration (CWE) instances in these code snippets. The results show that (1) 35.8% of Copilot generated code snippets contain CWEs, and those issues are spread across multiple languages, (2) the security weaknesses are diverse and related to 42 different CWEs, in which CWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values, and CWE-703: Improper Check or Handling of Exceptional Conditions occurred the most frequently, and (3) among the 42 CWEs identified, 11 of those belong to the currently recognized 2022 CWE Top-25. Our findings confirm that developers should be careful when adding code generated by Copilot (and similar AI code generation tools) and should also run appropriate security checks as they accept the suggested code. It also shows that practitioners should cultivate corresponding security awareness and skills.
Code Generation, Security Weaknesses, CWEs, GitHub Copilot
$c$s: Software and its engineering Software development techniquesccs: Security and privacy Software security engineering
@@ 1. Introduction
Code generation tools aim to automatically generate functional code based on prompts, which can include text descriptions (comments), code (such as function signatures, expressions, variable names, etc.), or a combination of text and code (Sarkar et al., 2022). After writing an initial code or comment, developers can rely on code generation tools to complete the remaining code. This approach can save development time and accelerate the software development process. Automated code generation tools have always been an active research discussion topic (Siddiq and Santos, 2022; Li et al., 2022b). Some of the earliest work can be traced back to the 1960s when Waldinger and Lee proposed a program synthesizer called PROW, which automatically generated LISP programs based on specifications provided by users in the form of predicate calculus (Waldinger and Lee, 1969). As computer languages continued to evolve, more and more programming languages began to support meta-programming, making automated code generation technology more efficient and flexible. In recent years, the rapid development of artificial intelligence technologies, particularly in the form of machine learning and deep learning models, has accelerated the development of code generation technologies.
@@
Recent advancements in code generation came with the emergence of Large Language Models (LLMs). LLMs are deep learning models trained on a large code/text corpus with powerful language understanding capabilities that can be used for tasks such as natural language generation, text classification, and question-answering systems (Chen et al., 2021). Compared to previous deep learning methods, the latest developments in LLMs, such as Generative Pre-trained Transformer (GPT) models, have opened up new opportunities to address the limitations of existing automated code generation technology (Mozannar et al., 2022). Currently, code generation tools based on LLMs have also been widely applied, such as Codex by OpenAI (OpenAI, [n. d.]), AlphaCode by DeepMind (Li et al., 2022a), and CodeWhisperer by Amazon (Becker et al., 2022).
@@
These models are trained on billions of public open-source lines of code, which includes public code with unsafe coding patterns (He and Vechev, 2023). Therefore, code generation tools based on such models can pose security risks, and the code they generate may also have security weaknesses. For example, GitHub Copilot may produce some insecure code, as its underlying model Codex is pre-trained on untrusted data from GitHub (Brown et al., 2020), which is known to contain buggy programs (Rokon et al., 2020). In addition, the code with vulnerabilities generated by these code generation tools may continue to be used to train the model, thus further generating code with vulnerabilities, leading to a vicious cycle. Previous research has studied code generation tools, with more focus on the correctness of the results (Lertbanjongngam et al., 2022; Wong et al., 2022; Dakhel et al., 2022; Pudari and Ernst, 2023), and relatively less attention has been paid to security aspects (Pearce et al., 2022; Siddiq et al., 2022; Perry et al., 2022). To the best of our knowledge, potential security weaknesses in practical scenarios have not been fully considered and addressed in previous work, and GitHub Copilot clarifies that “the users of Copilot are responsible for ensuring the security and quality of their code (GitHub, [n. d.])”. GitHub also provides tools such as CodeQL to help developers scan the security issues in their code.
@@
To this end, we conducted an empirical study on the security weaknesses of generated code by GitHub Copilot, which is available on GitHub. We chose Copilot as our research subject because it is a commercial instance of AI-assisted programming and has gained much attention and popularity among developers since its launch in 2021. The security weaknesses of code generated by Copilot have also gained attention in the research and practice community. Furthermore, thousands of developers in the GitHub community have shared their experiences of using Copilot in real-world systems (Dakhel et al., 2022). We collected code generated by Copilot that has been used in projects on GitHub and analyzed the security of the generated code through the lens of a real-world production environment. Then, we used static analysis tools to perform security analysis on the collected code snippets and classified the security weaknesses in the code snippets using the Common Weakness Enumeration (CWE).
@@
Our findings show that: (1) 35.8% of Copilot generated code snippets have security weaknesses, and security weaknesses arise regardless of the programming language used; (2) the security weaknesses are diverse and related to 42 different CWEs, in whichCWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values and CWE-703: Improper Check or Handling of Exceptional Conditions are the most frequently occurred; and (3) among the 42 CWEs identified, 11 CWEs belong to the currently recognized 2022 CWE Top-25.
@@
The contributions of this work: (1) We curated a dataset of code snippets generated by Copilot that has been used in projects on GitHub (a curated data (Fu et al., 2023) is made available online for future research in this area.) and conducted security checks on them, which can to some extent reflect the frequency of security weaknesses encountered by developers when using Copilot to generate code in actual coding; (2) We extensively checked all possible CWEs in the code snippets and analyzed them. This can help developers understand the common CWEs caused by using Copilot to generate code in actual coding and how to accept the code suggestions provided by Copilot safely.
@@
The rest of this paper is structured as follows: Section 2 presents the related work. Section 3 presents the research questions and the research design of this study. Section 4 presents our study results, which are further discussed in Section 5. The potential threats to validity are clarified in Section 6. Section 7 concludes this work with future work directions.
@@ 2.1. AI-assisted Code Generation Tools
With the rise of code generation tools integrated with IDEs, many studies have evaluated these code generation systems based on transformer models to better understand their effectiveness in real-world scenarios. Previous research mainly focused on whether the code generated by these tools can meet users’ functional requirements. Yetistiren et al. (Yetistiren et al., 2022) evaluated the effectiveness, correctness, and efficiency of the code generated by GitHub Copilot, and the results showed that GitHub Copilot could generate valid code with a success rate of 91.5%, making it a promising tool. Sobania et al. (Sobania et al., 2022) evaluated the correctness of the code generated by GitHub Copilot and compared the tool with an automatic program generator with a Genetic Programming (GP) architecture. They concluded there was no significant difference between the two methods on benchmark problems. Nguyen and Nadi (Nguyen and Nadi, 2022) conducted an empirical study using 33 LeetCode problems and created queries for Copilot in four different programming languages. They evaluated the correctness and comprehensibility of the code suggested by Copilot by running tests provided by LeetCode. They found that Copilot’s suggestions have lower complexity. Burak et al. (Yetiştiren et al., 2023) evaluated the code quality of AI-assisted code generation tools (GitHub Copilot, Amazon CodeWhisperer, and ChatGPT). They compared the improvements between the latest and older versions of Copilot and CodeWhisperer and found that the quality of generated code had improved.
In recent years, researchers have also started to focus on the experience of developers when using AI-assisted code generation tools and how the tools can improve productivity by observing their behavior. Vaithilingam et al. (Vaithilingam et al., 2022) studied how programmers use and perceive Copilot, and they found that while Copilot may not necessarily improve task completion time or success rate, it often provides a useful starting point. They also noted that participants faced difficulties in understanding, editing, and debugging the code snippets generated by Copilot. Barke et al. (Barke et al., 2022) presented the first theoretical analysis of how programmers interact with Copilot based on the observations of 20 participants. Sila et al. (Lertbanjongngam et al., 2022) conducted an empirical study on AlphaCode, identifying similarities and performance differences between code generated by code generation tools and code written by human developers. They argued that software developers should check the generated code for potentially problematic code that could introduce performance weaknesses.
These works above conducted relatively extensive evaluations of code generation tools in terms of correctness, effectiveness, and robustness. However, there is still room for improvement regarding its security, as detailed in the following section.
@@ 2.2. Security of Code Generation Techniques and LLMs
Code security is an issue that cannot be ignored in the software development process. Recent work has primarily focused on evaluating the security of the code generation tools and the security of the LLMs that these tools are based on.
@@
Pearce et al. (Pearce et al., 2022) first evaluated the security of GitHub Copilot in generating programs by identifying known weaknesses in the suggested code. The authors prompted Copilot to generate code for 89 cybersecurity scenarios and evaluated the weaknesses in the generated code. They found that 40% of the suggestions in the relevant context contained security-related bugs (i.e., CWE classification from MITRE (The MITRE Corporation, [n. d.])). Siddiq et al. (Siddiq et al., 2022) conducted a large-scale empirical study on code smells in the training set of a transformer-based Python code generation model and investigated the impact of these harmful patterns on the generated code. They observed that Copilot introduces 18 code smells, including non-standard coding patterns and two security smells (i.e., code patterns that often lead to security defects). Khoury et al. (Khoury et al., 2023) studied the security of the source code generated by the ChatGPT chatbot based on LLMs, and they found that ChatGPT was aware of potential weaknesses but still frequently generated some non-robust code.
@@
Several researchers also compared the situation where code generation tools produce insecure code with that of human developers. Sandoval et al. (Sandoval et al., 2022) conducted a security-driven user study, and their results showed that the rate at which AI-assisted user programming produced critical security errors was no more than 10% of the control group, indicating that the use of LLMs does not introduce new security risks. Asare et al. (Asare et al., 2023) conducted a comparative empirical analysis of these tools and language models from a security perspective and investigated whether Copilot is as bad as humans in generating insecure code. They found that while Copilot performs differently across vulnerability types, it is not as bad as human developers when it comes to introducing vulnerabilities in code. In addition, researchers have also constructed datasets to test the security of these tools. Tony et al. (Tony et al., 2023) proposed LLMSecEval, a dataset containing 150 natural language prompts that can be used to evaluate the security performance of LLMs. Siddiq et al. (Siddiq and Santos, 2022) provided a dataset, SecurityEval, for testing whether a code generation model has weaknesses. The dataset contains 130 Python code samples.
@@
Unlike the works above, we studied the security weaknesses exhibited by code generation tools in a real-world production environment (i.e., GitHub). We collected code snippets from GitHub generated by developers using Copilot in daily production as a source of research data, whereas in the Pearce et al. (Pearce et al., 2022) study, the research data came from code generated by the authors using Copilot based on the natural language prompts related to high-risk network security weaknesses. In addition to this, Pearce et al. configured CodeQL only to examine CWEs targeted by security weaknesses associated with the prompted scenarios. In contrast, we used various static analysis tools to examine all types of CWEs and analyze them extensively. Our research results may help developers understand what common CWEs are prone to result from using Copilot to generate code in real coding.
@@ 2.3. Security Static Analysis
Vulnerabilities detection is critical to improve software security and ensure quality  (Iannone et al., 2022). There are two used methods for vulnerability detection in source code: via static and dynamic code analysis. Dynamic analysis techniques are more sound and precise but lack coverage (Ernst, 2003). On the other hand, static analysis is less precise but offers greater coverage and allows to analyze programs without the need to execute them (Sui et al., 2020). Static analysis has been widely used to find security issues in code, given it is cheaper to run and can conduct whole program analyses without the need to execute the program (Dunlap et al., 2023). OWASP (OWASP, [n. d.]) provides a list of commonly used static analysis tools. This includes tools like CodeQL: a general-purpose automatic scanning tool, FindBugs: a tool for Java programs, ESLint: a tool for JavaScript programs, Bandit: a tool for Python programs, and GoSec: a tool for Go programs. Such tools have been widely used in previous security analysis research  (Tómasdóttir et al., 2018; Pearce et al., 2022; Siddiq et al., 2022).
Kaur et al. (Kaur and Nayyar, 2020) compared static analysis tools for vulnerability detection in scanning C/C++ and Java source code. Tomasdottir et al. (Tómasdóttir et al., 2018) conducted an empirical study on ESLint, the most commonly used JavaScript static analysis tool among developers. Pearce et al. (Pearce et al., 2022) used CodeQL for security weakness scanning of generated Python and C++ code. Siddiq et al. (Siddiq and Santos, 2022) used Bandit to check Python code generated using a test dataset.
These static analysis tools support different analysis algorithms and techniques. By using multiple tools for analysis, potential weaknesses in the code can be discovered from different perspectives and levels, avoiding omissions and improving the accuracy of the analysis. Our study first used CodeQL to scan the collected code snippets. CodeQL is an open-source tool that supports multiple languages, including Java, JavaScript, C++, C#, and Python. It can find weaknesses in a codebase based on a set of known weaknesses/rules. In addition, to obtain more comprehensive scan results, we supplemented the scan of code in different languages with static analysis tools (i.e., Cppcheck and Bandi) tailored to specific languages.

### v2

@@ Abstract.
Modern code generation tools, utilizing AI models like Large Language Models (LLMs), have gained popularity for producing functional code. However, their usage presents security challenges, often resulting in insecure code merging into the code base. Evaluating the quality of generated code, especially its security, is crucial. While prior research explored various aspects of code generation, the focus on security has been limited, mostly examining code produced in controlled environments rather than real-world scenarios. To address this gap, we conducted an empirical study, analyzing code snippets generated by GitHub Copilot from GitHub projects. Our analysis identified 452 snippets generated by Copilot, revealing a high likelihood of security issues, with 32.8% of Python and 24.5% of JavaScript snippets affected. These issues span 38 different Common Weakness Enumeration (CWE) categories, including significant ones like CWE-330: Use of Insufficiently Random Values, CWE-78: OS Command Injection, and CWE-94: Improper Control of Generation of Code. Notably, eight CWEs are among the 2023 CWE Top-25, highlighting their severity. Our findings confirm that developers should be careful when adding code generated by Copilot and should also run appropriate security checks as they accept the suggested code. It also shows that practitioners should cultivate corresponding security awareness and skills.
Code Generation, Security Weakness, CWE, GitHub Copilot, GitHub Project
journal: TOSEMccs: Software and its engineering Software development techniquesccs: Security and privacy Software security engineering
@@ 1. Introduction
Code generation tools aim to automatically generate functional code based on prompts, which can include text descriptions (comments), code (such as function signatures, expressions, variable names, etc.), or a combination of text and code (Sarkar et al., 2022). After writing an initial code or comment, developers can rely on code generation tools to complete the remaining code. This approach can save development time and accelerate the software development process. Automated code generation tools have always been a topic of active research (Siddiq and Santos, 2022; Li et al., 2022b). Some of the earliest work can be traced back to the 1960s, when Waldinger and Lee proposed a program synthesizer called PROW, which automatically generated LISP programs based on specifications provided by users in the form of a predicate calculus (Waldinger and Lee, 1969). As computer languages continued to evolve, more and more programming languages began to support meta-programming, making automated code generation technology more efficient and flexible. In recent years, the rapid development of artificial intelligence technologies, particularly machine learning and deep learning models, has accelerated the development of code generation technologies.
@@
Recent advancements in code generation came with the emergence of Large Language Models (LLMs). LLMs are deep learning models trained on a large code/text corpus with powerful language understanding capabilities that can be used for tasks such as natural language generation, text classification, and question-answer systems (Chen et al., 2021). Compared to previous deep learning methods, the latest developments in LLMs, such as Generative Pre-trained Transformer (GPT) models, have opened up new opportunities to address the limitations of existing automated code generation technology (Mozannar et al., 2022). Currently, LLM-based code generation tools have also been widely applied, such as Codex by OpenAI (OpenAI, [n. d.]), AlphaCode by DeepMind (Li et al., 2022a), and CodeWhisperer by Amazon (Becker et al., 2022).
These models are trained on billions of public open-source lines of code, which includes public code with unsafe coding patterns (He and Vechev, 2023). Therefore, code generation tools based on such models can pose code quality issues (Liu et al., 2024), and the code generated by these tools may also suffer from security weaknesses (She et al., 2023). For example, GitHub Copilot may produce some insecure code, as its underlying Codex model is pre-trained on untrusted data from GitHub (Brown et al., 2020), which is known to contain buggy programs (Rokon et al., 2020). According to the developer security company Snyk, GitHub Copilot may also replicate existing security issues in code to suggest insecure code when the user’s existing codebase contains security issues (Paul Krill, 2024).
@@
In addition, the code with vulnerabilities generated by these code-generation tools may continue to be used to train the model, thus further generating code with vulnerabilities, leading to a vicious cycle. Previous research has studied code generation tools, with more focus on the correctness of the results (Lertbanjongngam et al., 2022; Wong et al., 2022; Dakhel et al., 2022; Pudari and Ernst, 2023), and relatively less attention has been paid to security aspects (Pearce et al., 2022; Siddiq et al., 2022; Perry et al., 2022). To the best of our knowledge, potential security weaknesses in practical scenarios have not been fully considered and addressed in previous work, and GitHub Copilot clarifies that “the users of Copilot are responsible for ensuring the security and quality of their code” (GitHub, 2023b). Code generation algorithms of Copilot are incentivized to suggest code to be accepted rather than other code qualities, e.g., easy to read and understand, which has an adverse impact for the code quality generated by Copilot (Harding and Kloster, 2024). GitHub also provides tools such as CodeQL to help developers scan for security issues in their code.
@@
To this end, we conducted an empirical study on the security weaknesses of the generated code by GitHub Copilot, which is available on GitHub. We chose Copilot as our research subject because it is a commercial instance of AI-assisted programming and has gained much attention and popularity among developers since its launch in 2021. The security weaknesses of code generated by Copilot have also gained attention in the research and practice community. Furthermore, thousands of developers in the GitHub community have shared their experiences of using Copilot in real-world systems (Dakhel et al., 2022). We collected the code generated by Copilot that has been used in projects on GitHub and analyzed the security of the generated code through the lens of a real-world production environment. Then, we used static analysis tools to perform security analysis on the collected code snippets and classified the security weaknesses in the code snippets using the Common Weakness Enumeration (CWE).
@@
Our findings show that: (1) 29.6% of Copilot-generated code snippets have security weaknesses; (2) the security weaknesses are diverse and related to 38 different CWEs, in whichCWE-330: Use of Insufficiently Random Values, CWE-78: OS Command Injection and CWE-94: Improper Control of Generation of Code (’Code Injection’) are the most frequently occurred; and (3) among the 38 CWEs identified, eight CWEs belong to the currently recognized 2023 CWE Top-25. Six CWEs belong to Stubborn Weaknesses in the CWE Top 25.
@@
The contributions of this work: (1) We curated a dataset of code snippets generated by Copilot that has been used in projects on GitHub (a curated data (Fu et al., 2024) is made available online ) and conducted security checks on them, which can to some extent reflect the frequency of security weaknesses encountered by developers when using Copilot to generate code in actual coding. In addition to this, we also categorized the application areas of these code snippets; (2) We extensively checked all possible CWEs in the code snippets and analyzed them. This can help developers understand the common CWEs caused by using Copilot to generate code in actual coding and how to safely accept the code suggestions provided by Copilot.
@@
The rest of this paper is structured as follows: Section 2 presents the related work. Section 3 presents the research questions and the research design of this study. Section 4 presents the results of our study, which are further discussed in Section 5. The potential threats to validity are clarified in Section 6. Section 7 concludes this work with future work directions.
@@ 2. Related Work
In this section, we present the related work in three aspects, i.e., AI-assisted code generation tools (Section 2.1), security of code generation techniques and LLMs (Section 2.2), and static analysis tools for scanning security weaknesses (Section 2.3).
@@ 2.1. AI-assisted Code Generation Tools
With the rise of code generation tools integrated with IDEs, many studies have evaluated these systems based on transformer models to better understand their effectiveness in real-world scenarios. Previous research mainly focused on whether the code generated by these tools can meet users’ functional requirements. Yetistiren et al. (Yetistiren et al., 2022) evaluated the effectiveness, correctness, and efficiency of the code generated by GitHub Copilot, and the results showed that GitHub Copilot could generate valid code with a success rate of 91.5%, making it a promising tool. Sobania et al. (Sobania et al., 2022) evaluated the correctness of the code generated by GitHub Copilot and compared the tool with an automatic program generator with a Genetic Programming (GP) architecture. They concluded there was no significant difference between the two methods on benchmark problems. Nguyen and Nadi (Nguyen and Nadi, 2022) conducted an empirical study using 33 LeetCode problems and created queries for Copilot in four different programming languages. They evaluated the correctness and comprehensibility of the code suggested by Copilot by running tests provided by LeetCode. They found that Copilot’s suggestions have lower complexity. Burak et al. (Yetiştiren et al., 2023) evaluated the code quality of AI-assisted code generation tools (GitHub Copilot, Amazon CodeWhisperer, and ChatGPT). They compared the improvements between the latest and older versions of Copilot and CodeWhisperer and found that the quality of generated code had improved.
In recent years, researchers have also started to focus on the experience of developers when using AI-assisted code generation tools and how the tools can improve productivity by observing their behavior. Vaithilingam et al. (Vaithilingam et al., 2022) studied how programmers use and perceive Copilot and found that although Copilot may not necessarily improve task completion time or success rate, it often provides a useful starting point. They also noted that the participants had difficulties understanding, editing, and debugging the code snippets generated by Copilot. Barke et al. (Barke et al., 2022) presented the first theoretical analysis of how programmers interact with Copilot based on the observations of 20 participants. Sila et al. (Lertbanjongngam et al., 2022) conducted an empirical study on AlphaCode, identifying similarities and performance differences between code generated by code generation tools and code written by human developers. They argued that software developers should check the generated code for potentially problematic code that could introduce performance weaknesses.
The studies presented above have conducted a relatively extensive evaluation of code-generation tools regarding correctness, effectiveness, and robustness. However, its security still has room for improvement, as detailed below.
@@ 2.2. Security of Code Generation Techniques and LLMs
Code security is an issue that cannot be ignored in the software development process. Recent work has primarily focused on evaluating the security of the code generation tools and the security of the LLMs based on these tools.
@@
Pearce et al. (Pearce et al., 2022) first evaluated the security of GitHub Copilot in generating programs by identifying known weaknesses in the suggested code. The authors prompted Copilot to generate code for 89 cybersecurity scenarios and evaluated the weaknesses in the generated code. They found that 40% of the suggestions in the relevant context contained security-related bugs (i.e., CWE classification from MITRE (The MITRE Corporation, 2023b)). Siddiq et al. (Siddiq et al., 2022) conducted a large-scale empirical study on code smells in the training set of a transformer-based Python code generation model and investigated the impact of these harmful patterns on the generated code. They observed that Copilot introduces 18 code smells, including non-standard coding patterns and two security smells (i.e., code patterns that often lead to security defects). Khoury et al. (Khoury et al., 2023) studied the security of the source code generated by the ChatGPT chatbot based on LLMs, and found that ChatGPT was aware of potential weaknesses but still frequently generated some non-robust code. Elgedawy et al. (Elgedawy et al., 2024) compared the capabilities of four code generation models using nine code generation tasks. They collected 61 code outputs and studied their security. The results revealed that the code generated by different LLMs exhibited disparate levels of security robustness.
@@
Several researchers also compared the situation in which code generation tools produce insecure code with that of human developers. Sandoval et al. (Sandoval et al., 2022) conducted a security-driven user study, and their results showed that the rate at which AI-assisted user programming produced critical security errors was no more than 10% of the control group, indicating that the use of LLMs does not introduce new security risks. Asare et al. (Asare et al., 2023) conducted a comparative empirical analysis of these tools and language models from a security perspective and investigated whether Copilot is as bad as humans in generating insecure code. They found that while Copilot performs differently across vulnerability types, it is not as bad as human developers when introducing code vulnerabilities. In addition, researchers have also constructed datasets to test the security of these tools. Tony et al. (Tony et al., 2023) proposed LLMSecEval, a dataset containing 150 natural language prompts that can be used to evaluate the security performance of LLMs. Siddiq et al. (Siddiq and Santos, 2022) provided a dataset, SecurityEval, for testing whether a code generation model has weaknesses. The dataset contains 130 Python code samples.
@@
Different from prior work, we studied the security weaknesses exhibited by code generation tools in a real-world production environment (i.e., GitHub). We collected code snippets from GitHub generated by developers using Copilot in daily production as a source of research data, whereas in the Pearce et al. (Pearce et al., 2022) study, the research data came from code generated by the authors using Copilot based on natural language prompts related to high-risk network security weaknesses. Additionally, Pearce et al. configured CodeQL only to examine CWEs targeted by security weaknesses associated with the provided scenarios. In contrast, we used various static analysis tools to examine all types of CWEs and analyze them extensively. Our research results may help developers understand what common CWEs are prone to result from using Copilot to generate code in real coding.
@@ 2.3. Security Static Analysis
Vulnerabilities detection is critical to improve software security and ensure quality  (Iannone et al., 2022). Vulnerability can be checked either statically or dynamically. Dynamic analysis techniques are more sound and precise but lack coverage (Ernst, 2003). On the other hand, static analysis is less precise but offers greater coverage and allows users to analyze programs without the need to execute them (Sui et al., 2020). OWASP (OWASP, [n. d.]) provides a list of commonly used static analysis tools for security analysis. This includes tools like CodeQL, a general-purpose automatic scanning tool; FindBugs, a tool for Java programs; ESLint, a tool for JavaScript programs; Bandit, a tool for Python programs; and GoSec, a tool for Go programs. Such tools have been widely used in previous security analysis research  (Tómasdóttir et al., 2018; Pearce et al., 2022; Siddiq et al., 2022).
Kaur et al. (Kaur and Nayyar, 2020) compared static analysis tools for vulnerability detection in scanning C/C++ and Java source code. Tomasdottir et al. (Tómasdóttir et al., 2018) conducted an empirical study on ESLint, the most commonly used JavaScript static analysis tool among developers. Pearce et al. (Pearce et al., 2022) used CodeQL to scan security weaknesses in the generated Python and C++ code. Siddiq et al. (Siddiq and Santos, 2022) used Bandit to check Python code generated using a test dataset. Lisa et al. (Do et al., 2020) reported on users’ goals, motivations, and strategies when using static analysis tools.
These static analysis tools support different analysis algorithms and techniques. By using multiple tools for analysis, potential weaknesses in the code can be discovered from different perspectives and levels, avoiding omissions and improving the accuracy of the analysis. Our study first used CodeQL to scan the collected code snippets. CodeQL is an open-source tool that supports multiple languages, including Java, JavaScript, C++, C#, and Python. It can find weaknesses in a codebase based on known weaknesses/rules. In addition, to obtain more comprehensive scan results, we supplemented the scan of code in different languages with static analysis tools (i.e., Cppcheck and Bandi) tailored to specific languages.